OpenStudy (anonymous):

Use the quadratic formula to solve -4x2 + 4x + 80 = 0

OpenStudy (anonymous):

quadratic formula is \[(-b \pm \sqrt{b ^{2}-4ac})/2a\] Let a=-4, b=4, c=80 \[(-4 \pm \sqrt{16-4(-4)(80)})/ 2a\] We can see from the discriminant that this will have no real roots (the discriminant is the part under the radical, and it is very clearly negative. Since we can't take the square root of a negative in the real domain, we must use complex numbers). Would you like help solving for the complex numbers? I can help you if you need.